# Key-value memory in the brain

### v1

@@ 4.1 Retrieval interference, not erasure, is the principal limiting factor in memory performance
A number of attempts have been made to estimate the storage capacity of the human brain (summarized in Dudai,, 1997). Depending on different assumptions about numbers of neurons and connectivity, estimates can range from $10^{7}$ to $10^{15}$ bits. The total number of bits arriving from sensory systems with minimal compression are estimated to range from $10^{13}$ to $10^{17}$. These rough numbers suggest that with adequate compression, storage capacity may not be the strongest constraint on memory performance.
@@
Perhaps more persuasively than these theoretical arguments, we can make the case based on observations about behavior. If memory storage has reached its capacity limit, then it is impossible to store new information without removing some old information. This implies that old information should become permanently inacessible at some point. In contrast, studies of human memory demonstrate that memories can be stored over decades, despite being rarely rehearsed (Bahrick et al.,, 1975; Bahrick and Hall,, 1991; Conway et al.,, 1991; Maxcey et al.,, 2021).
@@
One might try to explain these findings by arguing that memory storage has not reached the capacity limit, but then it would be very challenging to explain forgetting over the much shorter intervals studied in experiments on list memory. When presented with a list of random words, the proportion of recalled words declines with list length (typically in the range of 5 to 20 items). If this was due to removal of items from memory, then one would expect catastrophic forgetting over intervals of decades. Furthermore, a series of experiments by Shiffrin, (1970) demonstrated that, when presented with a sequence of lists and then asked to recall the list before the most recently presented list, performance depended not on the length of the most recent list but only on the length of the list being recalled. This suggests that forgetting is not due to displacement of old items by items from the last list. The limiting factor is retrieval interference from other items on the same list.
@@
Using word-location pairs, Berens et al., (2020) separately estimated memory accessibility (whether or not a location is recalled at all given a word cue) and precision (the variance of angular errors). They found that accessibility, but not precision, declined as a function of the retention interval. Similar results have been reported using memory for real-world events (Diamond et al.,, 2020). Thus, memories do not melt into oblivion, but rather disappear from view. When they come back into view, they are as sharp as they were before they disappeared.
@@
Supposedly lost memories can be found when appropriate retrieval cues are used (Wagenaar,, 1986), when the number of retrieval attempts increases (Buschke,, 1974; Roediger and Thorpe,, 1978), or even spontaneously after a sufficiently long delay (Payne,, 1987). This holds not only for standard memory tasks in healthy subjects but also for retrograde amnesia induced experimentally or by neurological damage. Spontaneous “shrinkage” of amnesia is a common clinical observation following brain trauma (Kapur,, 1999), presumably due to the restoration of memory access. In laboratory studies, amnesia can be induced experimentally in a range of ways, such as electroconvulsive shock, hypothermia, protein synthesis inhibition, and lesion or inactivation of the hippocampus, with recovery also induced in a range of ways (Lewis et al.,, 1968; Riccio and Richardson,, 1984; Miller,, 2021). For example, protein synthesis inhibition following classical conditioning typically eliminates conditioned responding on long-term memory tests. However, delayed testing sometimes reveals recovery of performance (Flexner et al.,, 1966; Serota,, 1971; Squire and Barondes,, 1972). It is even possible to restore performance using the amnestic agent itself (Bradley and Galal,, 1988; Briggs and Olson,, 2013; Gisquet-Verrier et al.,, 2015).
@@
These observations about amnesia mirror well-known phenomena in classical conditioning. Extinction of a previously conditioned stimulus (i.e., presenting the stimulus in the absence of the unconditioned stimulus) causes a decline in conditioned responding, eventually reaching baseline levels. This decline is transient: conditioned responding can return spontaneously (Pavlov,, 1927), or can be induced by a single “reminder” of the unconditioned stimulus (Rescorla and Heth,, 1975).
@@
In summary, considerable evidence suggests that failures of remembering primarily arise from failures of retrieval, typically due to interference from other memories. Memories thought to be lost can later be found under the right conditions. We will explore this idea computationally when we discuss model simulations (Sec. 5.2).
@@ 4.2 Distinct representations of keys and values
The influential “Complementary Learning Systems” framework holds that there is a division of labor between the hippocampus and neocortex (McClelland et al.,, 1995; O’Reilly and Norman,, 2002), with the hippocampus specialized for episodic memory (remembering events that occurred in a specific spatiotemporal context) and a set of neocortical areas (sometimes referred to as “association cortex”) that are specialized for semantic memory (remembering regularities that generalize across episodes). This framework has also influenced the design of artificial intelligence systems (Kumaran et al.,, 2016). In this section, we will argue that this framework can be understood in terms of key-value memory.
@@
Rather than thinking about the hippocampus as directly storing memory content, we can alternatively conceptualize it as storing keys and matching queries to keys, which address memory content stored in neocortex. This view emphasizes the point that episodic memories need to be bound to semantic content—otherwise, they’re essentially empty vessels, as in cases of semantic demantia, where degeneration of anterior temporal lobe and prefrontal areas produces profound semantic impairments despite relatively intact recognition memory for recent experiences (Graham et al.,, 1999).
@@
If the hippocampus provides the keys necessary for activating neocortical values, then we would expect a causal interplay between the two. Indeed, there is empirical evidence for the following facts: (i) cortical encoding-related activity is reinstated at the time of memory retrieval; (ii) cortical reinstatement depends on the hippocampus; and (iii) the reinstatement is necessary for memory retrieval (Staresina et al.,, 2012; Bosch et al.,, 2014; Tanaka et al.,, 2014; Danker et al.,, 2017; Pacheco Estefan et al.,, 2019; Hebscher et al.,, 2021).
@@
Another line of evidence comes from studies of generalization. In the absence of the hippocampus, neocortical values cannot be accessed in a targeted way, leading to overgeneralization. A study by Winocur et al., (2009) offers a good example. When trained to anticipate a shock in Context A, rats specifically freeze in Context A but not in Context B when tested a day later. When tested a week later, rats show a generalization effect (loss of context specificity), freezing in both contexts. This generalization might arise because of interference from memories acquired during the intervening time. Context specificity can be restored by “reminding” the rat of the original memory (briefly placing it back in Context A). This reminder effect can be interpreted as activation of the appropriate address given a highly specific cue. Importantly, the reminder effect disappears in hippocampal-lesioned rats, consistent with the claim that the hippocampus stores the keys necessary for targeting specific memories (see also Wiltgen et al.,, 2010, for converging evidence).
@@
The idea that the hippocampus stores keys was to a large extent anticipated by the hippocampal memory indexing theory of Teyler and DiScenna, (1986), which proposed that the hippocampus indexes memory content stored in neocortical areas. Since it was first proposed, new techniques have uncovered much more detailed support for the theory (Teyler and Rudy,, 2007; Goode et al.,, 2020). In particular, the advent of activity-dependent labeling and optogenetic manipulation have enabled the identification of “engram cells” in the hippocampus which are causally linked to specific memories (Liu et al.,, 2012; Ramirez et al.,, 2013). Goode et al., (2020) interpret hippocampal engrams as indices (in the sense of Teyler and DiScenna), linking together information stored in a distributed network of neocortical areas.
@@
Several brain-wide engram mapping studies in rodents have reported a collection of neocortical (as well as subcortical) areas that are active during both encoding and retrieval, and therefore qualify as engrams (Vetere et al.,, 2017; Roy et al.,, 2022). What makes the hippocampus special is its role as a hub region with high connectivity to neocortical regions (Battaglia et al.,, 2011). This allows the hippocampus to exert strong control over neocortical regions. In addition, hippocampal engrams are highly sparse (involving a small subset of hippocampal cells) and conjunctively tuned to multiple sensory and cognitive inputs; these features make hippocampal engrams particularly well-suited to encoding episodically unique memories. A recent study of food-caching birds (Chettih et al.,, 2024) provides a particularly dramatic demonstration: hippocampal ensembles encoded unique representations for over 100 cache sites (including multiple caches at the same location), which were reactivated upon memory retrieval.
@@
If hippocampal representations are optimized for discriminating between distinct episodes in a cue-dependent manner, then we should expect changes in these representations under different retrieval demands. For example, overlapping routes in spatial navigation need to be discriminated in order to avoid confusion. Chanales et al., (2017) showed that this situation produces repulsion of hippocampal representations specifically for overlapping locations (see also Wanjia et al.,, 2021). The repulsion effect emerges gradually over the course of learning, ultimately reversing the objective similarity relations between locations. The strength of the repulsion effect also correlated with behaviorally measured discrimination accuracy.
@@
Note that optimization for discriminability is only part of the story, since the hippocampus receives noisy inputs which may activate the wrong keys. The hippocampus needs to do some error correction / pattern completion in order to “clean up” the retrieved keys. This may be accomplished by attractor dynamics such that nearby inputs get mapped to a corresponding attractor (pattern completion), while more distant inputs are separated into distinct attractors (pattern separation). It remains to be seen if such attractor dynamics are implemented by the recurrent hippocampal-entorhinal loop (Chandra et al.,, 2023) (Box 1) or by the proposal that both occur within hippocampus, with pattern separation by dentate gyrus and pattern completion by area CA3 (Rolls,, 2013).
@@ 4.3 Values, but not keys, are available for recall
Many of us are familiar with the experience of having a memory at the “tip of the tongue”—stored in memory yet temporarily unrecallable (Brown and McNeill,, 1966; Brown,, 1991). Closely related is the “feeling of knowing”—a subjective judgment about the likelihood of subsequently recognizing items which are presently unrecallable. The typical study procedure is to present subjects with difficult general knowledge questions, elicit feelings of knowing for questions they cannot presently answer, and then subsequently test their ability to recognize correct answers to the questions. An important finding from this literature is that feelings of knowing predict (though not perfectly) subsequent recognition (Hart,, 1965; Freedman and Landauer,, 1966), indicating that people are able to judge whether some piece of information is stored in memory despite not being able to retrieve it. Similar results have been found with cued recall tests (Gruneberg and Monks,, 1974). Another clue comes from studies examining response times: Reder, (1987) found that people could report answerability for trivia questions faster than they could report the answers, again indicating that metamemory does not require retrieval of memory content. Furthermore, feelings of knowing (but not recall) can be spuriously increased by increasing cue familiarity (Reder and Ritter,, 1992; Schwartz and Metcalfe,, 1992), possibly due to increased key-query match without increased probability that the correct keys are matched.
These phenomena are broadly consistent with the hypothesis that key-query matching can be used to judge whether some information (e.g., the answer to a question) is stored in memory, without accessing the memory content (value) itself. This idea has appeared in the human long-term memory literature under various guises; for example, Koriat and Lieblich, (1977) discuss the tip-of-the-tongue phenomenon in terms of “pointers” (cues that specify a memory address without specifying memory content), whereas Morton et al., (1985) use “headings” to refer to essentially the same idea.
A pointer system has also been invoked to understand short-term memory: rather than storing transient copies of long-term memories, short-term memory might store pointers which refer to information in long-term memory (Ruchkin et al.,, 2003; Norris,, 2017). This may explain why people can detect changes even when they cannot report what exactly has changed (Ball and Busch,, 2015), analogous to the “butcher on the bus” phenomenon in long term memory (Mandler,, 1980), where people can recognize familiar items without being able to recollect any details about them. Both change detection without identification and familiarity without recollection might be accomplished using keys (pointers) that are content addressable but do not obligatorily activate their associated values (Chandra et al.,, 2023).
@@ 5.2 Forgetting as retrieval failure, and recovery by memory reactivation
Another distinctive property of the key-value memory we highlighted in Sec. 4.1 is that, as recall relies on successful retrieval, forgetting can be conceptualized as retrieval failure; that is, even when a memory trace (i.e., a key/value pair corresponding to an event) is stored in memory, forgetting can still occur when retrieval fails due to interference/inaccessibility. This also implies that, if we manage to fix the failure in the retrieval process, there is a hope for recovering the corresponding memory without having to relive the event itself. Here we present a simulation illustrating these phenomena using an artificial neural network that learns two tasks sequentially (the so-called continual learning scenario), and examine it through the lens of key-value memory formulation of Sec. 2.3. We also highlight how this simple experiment echoes neurobiological findings on optogenetic recovery of memory following retrograde amnesia (Ryan et al.,, 2015; Roy et al.,, 2017).
We train a simple feedforward neural network on two binary image classification tasks sequentially. Using two classic image datasets commonly used in deep learning, MNIST and FashionMNIST (LeCun et al.,, 1998; Xiao et al.,, 2017), we construct the two toy tasks that involve classifying images from the two first classes of MNIST (digit ‘0’ vs. ‘1’) and FashionMNIST (‘T-shirt’ vs. ‘Trouser’), respectively. The model has one hidden layer (i.e., it has two linear transformations: input-to-hidden, and hidden-to-output mappings with a rectifier linear activation function in-between). The input images are grayscale and their dimension is 28$\times$28; they are flattened to yield a 784-dimensional vector accepted by the input layer. We set the hidden-layer dimension to 64, and the model output size is 4 (for two times two-way classification tasks). The model is trained on the cross-entropy loss using the gradient descent algorithm; by applying the dual formulation described in Sec. 2.3, each linear layer in the model can be formalized as a key-value system keeping memory traces of the entire learning experience by storing layer inputs as keys and error signals as values for every learning step.
@@
Figure 3A shows the evolution of model performance on the test set of the two tasks as a function of training epochs. We observe that the model achieves $\sim$99test accuracy on Task 1 (MNIST) in the first 5 epochs corresponding to Task 1 learning, but this performance drops to $\sim$9after learning Task 2 (FashionMNIST) until the model achieves $\sim$95test accuracy on Task 2 (we deliberately train the model long enough to observe forgetting). This amnesic phenomenon, reminiscent of catastrophic forgetting in neural networks (McCloskey and Cohen,, 1989; Ratcliff,, 1990; French,, 1999), is intriguing given the key-value memory view of the model, in which the key/value memory pairs belonging to Task 1 remain part of the model parameters, explicitly and indefinitely. By using the same terminology as in neuroscience experiments (Ryan et al.,, 2015), these Task 1-related key/value memories became “silent” after Task 2 learning.
An interesting question is whether we can reactivate these silent key/value memories to recover the model’s performance on Task 1 without any retraining on Task 1, akin to how experimental neuroscientists successfully reactivated “silent engrams” through an optogenetic procedure (Roy et al.,, 2017). For this, we introduce a single positive scalar $\beta\geq 1$ (an “optogenetic strength”) to multiply/amplify the keys (or, equivalently in this model, the values) in all key-value pairs corresponding to Task 1 in the trained model. Figure 3B shows the performance of the model on Task 1 (MNIST) as a function of the optogenetic strength $\beta$ ($\beta=1$ corresponds to no intervention on the model). We observe that by simply increasing $\beta$ (i.e., by artificially “reactivating” the existing key-value memories corresponding to the model’s prior MNIST learning experiences), the model becomes capable of solving MNIST again without any retraining.

### v2

@@ 4.1 Retrieval interference, not erasure, is the principal limiting factor in memory performance
A number of attempts have been made to estimate the storage capacity of the human brain 68. Depending on different assumptions about numbers of neurons and connectivity, estimates can range from $10^{7}$ to $10^{15}$ bits. The total number of bits arriving from sensory systems with minimal compression are estimated to range from $10^{13}$ to $10^{17}$. These rough numbers suggest that with adequate compression, storage capacity may not be the strongest constraint on memory performance.
@@
Perhaps more persuasively than these theoretical arguments, we can make the case based on observations about behavior. If memory storage has reached its capacity limit, then it is impossible to store new information without removing some old information. This implies that old information should become permanently inacessible at some point. In contrast, studies of human memory demonstrate that memories can be stored over decades, despite being rarely rehearsed 69, 70, 71, 72.
@@
One might try to explain these findings by arguing that memory storage has not reached the capacity limit, but then it would be very challenging to explain forgetting over the much shorter intervals studied in experiments on list memory. When presented with a list of random words, the proportion of recalled words declines with list length (typically in the range of 5 to 20 items). If this was due to removal of items from memory, then one would expect catastrophic forgetting over intervals of decades. Furthermore, a series of experiments by Shiffrin 73 demonstrated that, when presented with a sequence of lists and then asked to recall the list before the most recently presented list, performance depended not on the length of the most recent list but only on the length of the list being recalled. This suggests that forgetting is not due to displacement of old items by items from the last list. The limiting factor is retrieval interference from other items on the same list.
@@
Using word-location pairs, Berens and colleagues 74 separately estimated memory accessibility (whether or not a location is recalled at all given a word cue) and precision (the variance of angular errors). They found that accessibility, but not precision, declined as a function of the retention interval. Similar results have been reported using memory for real-world events 75. Thus, memories do not melt into oblivion, but rather disappear from view. When they come back into view, they are as sharp as they were before they disappeared.
@@
Supposedly lost memories can be found when appropriate retrieval cues are used 76, when the number of retrieval attempts increases 77, 78, or even spontaneously after a sufficiently long delay 79. This holds not only for standard memory tasks in healthy subjects but also for retrograde amnesia induced experimentally or by neurological damage. Spontaneous “shrinkage” of amnesia is a common clinical observation following brain trauma 80, presumably due to the restoration of memory access. In laboratory studies, amnesia can be induced experimentally in a range of ways, such as electroconvulsive shock, hypothermia, protein synthesis inhibition, and lesion or inactivation of the hippocampus, with recovery also induced in a range of ways 81, 82, 4. For example, protein synthesis inhibition following classical conditioning typically eliminates conditioned responding on long-term memory tests. However, delayed testing sometimes reveals recovery of performance 83, 84, 85. It is even possible to restore performance using the amnestic agent itself 86, 87, 88.
@@
These observations about amnesia mirror well-known phenomena in classical conditioning. Extinction of a previously conditioned stimulus (i.e., presenting the stimulus in the absence of the unconditioned stimulus) causes a decline in conditioned responding, eventually reaching baseline levels. This decline is transient: conditioned responding can return spontaneously 89, or can be induced by a single “reminder” of the unconditioned stimulus 90.
@@
In summary, considerable evidence suggests that failures of remembering primarily arise from failures of retrieval, typically due to interference from other memories. Memories thought to be lost can later be found under the right conditions. We will explore this idea computationally when we discuss model simulations.
@@ 4.2 Distinct representations of keys and values
The influential “Complementary Learning Systems” framework holds that there is a division of labor between the hippocampus and neocortex 91, 92, with the hippocampus specialized for episodic memory (remembering events that occurred in a specific spatiotemporal context) and a set of neocortical areas (sometimes referred to as “association cortex”) that are specialized for semantic memory (remembering regularities that generalize across episodes). This framework has also influenced the design of artificial intelligence systems 93. In this section, we will argue that this framework can be understood in terms of key-value memory.
@@
Rather than thinking about the hippocampus as directly storing memory content, we can alternatively conceptualize it as storing keys and matching queries to keys, which address memory content stored in neocortex. This view emphasizes the point that episodic memories need to be bound to semantic content—otherwise, they’re essentially empty vessels, as in cases of semantic demantia, where degeneration of anterior temporal lobe and prefrontal areas produces profound semantic impairments despite relatively intact recognition memory for recent experiences 94.
@@
If the hippocampus provides the keys necessary for activating neocortical values, then we would expect a causal interplay between the two. Indeed, there is empirical evidence for the following facts: (i) cortical encoding-related activity is reinstated at the time of memory retrieval; (ii) cortical reinstatement depends on the hippocampus; and (iii) the reinstatement is necessary for memory retrieval 95, 96, 97, 98, 99, 100.
@@
Another line of evidence comes from studies of generalization. In the absence of the hippocampus, neocortical values cannot be accessed in a targeted way, leading to overgeneralization. A study by Winocur and colleagues 101 offers a good example. When trained to anticipate a shock in Context A, rats specifically freeze in Context A but not in Context B when tested a day later. When tested a week later, rats show a generalization effect (loss of context specificity), freezing in both contexts. This generalization might arise because of interference from memories acquired during the intervening time. Context specificity can be restored by “reminding” the rat of the original memory (briefly placing it back in Context A). This reminder effect can be interpreted as activation of the appropriate address given a highly specific cue. Importantly, the reminder effect disappears in hippocampal-lesioned rats, consistent with the claim that the hippocampus stores the keys necessary for targeting specific memories (see also 102 for converging evidence).
@@
The idea that the hippocampus stores keys was to a large extent anticipated by the hippocampal memory indexing theory 103, which proposed that the hippocampus indexes memory content stored in neocortical areas. Since it was first proposed, new techniques have uncovered much more detailed support for the theory 104, 105. In particular, the advent of activity-dependent labeling and optogenetic manipulation have enabled the identification of “engram cells” in the hippocampus which are causally linked to specific memories 106, 107. Goode and colleagues 105 interpret hippocampal engrams as indices (in the sense of Teyler and DiScenna), linking together information stored in a distributed network of neocortical areas.
@@
Several brain-wide engram mapping studies in rodents have reported a collection of neocortical (as well as subcortical) areas that are active during both encoding and retrieval, and therefore qualify as engrams 108, 109. What makes the hippocampus special is its role as a hub region with high connectivity to neocortical regions 110. This allows the hippocampus to exert strong control over neocortical regions. In addition, hippocampal engrams are highly sparse (involving a small subset of hippocampal cells) and conjunctively tuned to multiple sensory and cognitive inputs; these features make hippocampal engrams well-suited to encoding episodically unique memories. A recent study of food-caching birds 111 provides a particularly dramatic demonstration: hippocampal ensembles encoded unique representations for over 100 cache sites (including multiple caches at the same location), which were reactivated upon memory retrieval.
@@
If hippocampal representations are optimized for discriminating between distinct episodes in a cue-dependent manner, then we should expect changes in these representations under different retrieval demands. For example, overlapping routes in spatial navigation need to be discriminated in order to avoid confusion. Chanales and colleagues 112 showed that this situation produces repulsion of hippocampal representations specifically for overlapping locations (see also 113). The repulsion effect emerges gradually over the course of learning, ultimately reversing the objective similarity relations between locations. The strength of the repulsion effect also correlated with behaviorally measured discrimination accuracy.
@@
Note that optimization for discriminability is only part of the story, since the hippocampus receives noisy inputs which may activate the wrong keys. The hippocampus needs to do some error correction / pattern completion in order to “clean up” the retrieved keys. This may be accomplished by attractor dynamics such that nearby inputs get mapped to a corresponding attractor (pattern completion), while more distant inputs are separated into distinct attractors (pattern separation). It remains to be seen if such attractor dynamics are implemented by the recurrent hippocampal-entorhinal loop 55 (Box 1) or by the proposal that both occur within hippocampus, with pattern separation by dentate gyrus and pattern completion by area CA3 114.
@@ 4.3 Values, but not keys, are available for recall
Many of us are familiar with the experience of having a memory at the “tip of the tongue”—stored in memory yet temporarily unrecallable 115, 116. Closely related is the “feeling of knowing”—a subjective judgment about the likelihood of subsequently recognizing items which are presently unrecallable. The typical study procedure is to present subjects with difficult general knowledge questions, elicit feelings of knowing for questions they cannot presently answer, and then subsequently test their ability to recognize correct answers to the questions. An important finding from this literature is that feelings of knowing predict (though not perfectly) subsequent recognition 117, 118, indicating that people are able to judge whether some piece of information is stored in memory despite not being able to retrieve it. Similar results have been found with cued recall tests 119. Another clue comes from studies examining response times: Reder 120 found that people could report answerability for trivia questions faster than they could report the answers, again indicating that metamemory does not require retrieval of memory content. Furthermore, feelings of knowing (but not recall) can be spuriously increased by increasing cue familiarity 121, 122, possibly due to increased key-query match without increased probability that the correct keys are matched.
These phenomena are broadly consistent with the hypothesis that key-query matching can be used to judge whether some information (e.g., the answer to a question) is stored in memory, without accessing the memory content (value) itself. This idea has appeared in the human long-term memory literature under various guises; for example, Koriat 123 discuss the tip-of-the-tongue phenomenon in terms of “pointers” (cues that specify a memory address without specifying memory content), whereas Morton and colleagues 124 use “headings” to refer to essentially the same idea.
A pointer system has also been invoked to understand short-term memory: rather than storing transient copies of long-term memories, short-term memory might store pointers which refer to information in long-term memory 125, 126. This may explain why people can detect changes even when they cannot report what exactly has changed 127, analogous to the “butcher on the bus” phenomenon in long term memory 128, where people can recognize familiar items without being able to recollect any details about them. Both change detection without identification and familiarity without recollection might be accomplished using keys (pointers) that are content addressable but do not obligatorily activate their associated values 55.
@@ 5.2 Forgetting as retrieval failure, and recovery by memory reactivation
Another distinctive property of the key-value memory we highlighted earlier is that, as recall relies on successful retrieval, forgetting can be conceptualized as retrieval failure; that is, even when a memory trace (i.e., a key/value pair corresponding to an event) is stored in memory, forgetting can still occur when retrieval fails due to interference/inaccessibility. This also implies that, if we manage to fix the failure in the retrieval process, there is a hope for recovering the corresponding memory without having to relive the event itself. Here we present a simulation illustrating these phenomena using an artificial neural network that learns two tasks sequentially (the so-called continual learning scenario), and examine it through the lens of key-value memory. We also highlight how this simple experiment echoes neurobiological findings on optogenetic recovery of memory following retrograde amnesia 129, 130.
We train a simple feedforward neural network on two binary image classification tasks sequentially. Using two classic image datasets commonly used in deep learning, MNIST and FashionMNIST 131, 132, we construct the two toy tasks that involve classifying images from the two first classes of MNIST (digit ‘0’ vs. ‘1’) and FashionMNIST (‘T-shirt’ vs. ‘Trouser’), respectively. The model has one hidden layer (i.e., it has two linear transformations: input-to-hidden, and hidden-to-output mappings with a rectifier linear activation function in-between). The input images are grayscale and their dimension is 28$\times$28; they are flattened to yield a 784-dimensional vector accepted by the input layer. We set the hidden-layer dimension to 64, and the model output size is 4 (for two times two-way classification tasks). The model is trained on the cross-entropy loss using the gradient descent algorithm; by applying the dual formulation described earlier, each linear layer in the model can be formalized as a key-value system keeping memory traces of the entire learning experience by storing layer inputs as keys and error signals as values for every learning step.
@@
Figure 3A shows the evolution of model performance on the test set of the two tasks as a function of training epochs. We observe that the model achieves $\sim$99test accuracy on Task 1 (MNIST) in the first 5 epochs corresponding to Task 1 learning, but this performance drops to $\sim$9after learning Task 2 (FashionMNIST) until the model achieves $\sim$95test accuracy on Task 2 (we deliberately train the model long enough to observe forgetting). This amnesic phenomenon, reminiscent of catastrophic forgetting in neural networks 65, 133, 134, is intriguing given the key-value memory view of the model, in which the key/value memory pairs belonging to Task 1 remain part of the model parameters, explicitly and indefinitely. By using the same terminology as in neuroscience experiments 129, these Task 1-related key/value memories became “silent” after Task 2 learning.
An interesting question is whether we can reactivate these silent key/value memories to recover the model’s performance on Task 1 without any retraining on Task 1, akin to how experimental neuroscientists successfully reactivated “silent engrams” through an optogenetic procedure 130. For this, we introduce a single positive scalar $\beta\geq 1$ (an “optogenetic strength”) to multiply/amplify the keys (or, equivalently in this model, the values) in all key-value pairs corresponding to Task 1 in the trained model. Figure 3B shows the performance of the model on Task 1 (MNIST) as a function of the optogenetic strength $\beta$ ($\beta=1$ corresponds to no intervention on the model). We observe that by simply increasing $\beta$ (i.e., by artificially “reactivating” the existing key-value memories corresponding to the model’s prior MNIST learning experiences), the model becomes capable of solving MNIST again without any retraining.